OpenStudy (anonymous):

Solve Quadratic Formula : 2x^2-3x+1 .

OpenStudy (anonymous):

\[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

I Set It Up But After I Get The Answer , I'm Suppose Too Simplify But The Answer I Have , Not On The Answer Key . ( If I Did It Right . )

OpenStudy (anonymous):

x=1/1 and 1/2